# Active Collaborative Visual SLAM exploiting ORB Features

### v1

@@ I INTRODUCTION
Current research is increasingly focusing on Active Collaborative SLAM (AC-SLAM), which leverages multiple robots working in unison. This collaborative method offers significant advantages, such as faster terrain mapping and robust operation in dynamic scenarios. However, the deployment of multiple robots introduces challenges like coordination, resource management, and sensor data integration. Moreover, merging individual robot contributions into a unified map poses complex computational and algorithmic challenges. Here, we propose an implementation of a visual AC-SLAM algorithm using heterogeneous agents and extend the work in [1] to a multi-agent system, where a team of ground robots and aerial robots (UAVs) collaboratively map and explore the environment. To achieve this aim, we propose a method that exploits ORB feature information from different viewpoints of UAVs and ground robots into a common merged map, we detect frontiers in local and overlapped ground robot maps and propose a frontier filtering method that encourages the spread of agents and exploration of the environment. Additionally, we also present a re-localization scheme to guide robots to previously visited goal positions to favor loop closure.
@@ II-A Active Collaborative SLAM
[7] presents a decentralized method for a long-planning horizon of actions for exploration and maintains estimation uncertainties at a certain threshold. The active path planner uses a modified version of RRT* and an action is chosen that best minimizes the entropy change per distance traveled. The main advantage of this approach is that it maintains good pose estimation and encourages loop-closure trajectories. An interesting solution is given by a similar approach to the method proposed by [8] using a relative entropy (RE)-optimization method which integrates motion planning with robot localization and selects trajectories that minimize the localization error and associated uncertainty bound. A planning-cost function is computed, which includes the uncertainty in the state in addition to the state and control cost.
@@ II-B Frontiers-based exploration
Frontiers play a pivotal role in augmenting the precision of robot localization by enabling intelligent exploration and data acquisition strategies, effectively reducing uncertainty, and enhancing the map-building and localization processes. In [9] each frontier is segmented, a trajectory is planned for each segment, and the trajectory with the highest map-segment covariance is selected from the global-cost map. The work presented in [3] uses frontier exploration for autonomous exploration a utility function based on Shannon’s and Renyi entropy is used for the computation of the utility of paths. The method described by [10] uses a cost function that is somewhat similar to [4], which takes into consideration the discovery of the target area of a robot by another member of the swarm and switches from a frontier to a distance-based navigation function to guide the robot toward the goal frontier. Frontiers-based coverage approaches in [11] divide the perception task into a broad exploration layer and a detailed mapping layer, making use of heterogeneous robots to carry out the two tasks while solving a Fixed Start Open Traveling Salesman Problem (FSOTSP). Once a frontier has been identified, the robot can use path planning algorithms to reach it and maximize the exploration while minimizing its SLAM uncertainly. In [12] the authors present decentralized autonomous exploration method which takes into account frontiers detected by other agent and inter robot exchange of frontiers. This method uses distance and information gain matrices for frontier utility computation.
@@ III METHODOLOGY
While many research works have been focused on collaborative strategies for SLAM, or single-robot active-SLAM, only a few works have dealt with AC-SLAM. However, these approaches present common limitations: a) they have high computational costs associated with the number of frontiers processed. b) They fail to encourage the spread of robots into the environment by not taking into account the common frontier points from the agents. c) The uncertainty is quantified by a scalar mapping of the entire pose graph covariance matrix which may become very large, especially in landmark-based SLAM methods increasing the computational cost. d) In case of SLAM failure or high uncertainty, there is no re-localization method. Furthermore, they do not explicitly implement strategies for efficient management of frontiers to speed up map discovery and robot localization. In this work, we propose a visual AC-SLAM approach that deals with overcoming these limitations using a fleet of ground and aerial robots (UAVs) and exploiting visual features from different viewpoints of both heterogeneous robots. Our proposed method outlines a strategy aimed at reducing the number of frontiers for reward computation and distributing robots within the environment, thereby facilitating exploration and mapping. Furthermore, we also present a re-localization method quantifying the SLAM pose graph edge covariance matrix hereby allowing the robots to visit already visited goal positions for re-localization and uncertainty reduction for a better SLAM estimate. We leverage a combination of reward metrics, common map points, and merged map information gain to refine the goal selection.
@@
We expand [1] to a multi-robot AC-SLAM and incorporate the utility function from our previous work in [14] which considers frontier path entropy to exploit map uncertainty for utility computation. Utilizing the Intersection over Union (IoU) overlapping map information, the number of frontiers is significantly reduced for efficient frontier sharing and exploration. Utilizing edge D-Optimality as an uncertainty metric, we propose a re-localization scheme to encourage loop closure and reduce uncertainty. In Figure 1 we show the resultant architecture of the proposed system implemented in ROS Noetic using actionlib111http://wiki.ros.org/actionlib library for drone (UAV) and ground robots (Robot_0… Robot_N). Each robot/drone runs its own instance of ORBSLAM2 using onboard RGBD sensors. The resultant ORB [15] features are converted to a 3D voxel map (Octomap) [16] for efficient representation using Octo and Grid Mapper module222http://wiki.ros.org/octomap_server. For each of the ground robot and drone, we limit the maximum and minimum height of the Octomap to 0.5 and 1 meters () respectively, thus limiting the ground robots to ignore ORB features above 0.5 high and for drones to ignore them of less than 1. We then project this Octomap to a 2D Occupancy Grid (O.G) map and apply frontier detection methods using OpenCV and RRT-exploration from [1] to detect local frontiers in each robot’s local map. The MapMerger module merges the offline drone O.G maps with those of ground robots with world reference frame set to that of Robot_0. The IoU mapper computes the overlapped O.G map between two robots and detects frontiers and penalizes each robot’s local frontiers thus reducing them. Section III-A elaborates this process. Once the frontiers have been filtered they are passed to the frontier management and reward computation server as discussed in Section III-B which performs further reduction and updates the rewards before finally assigning goal frontiers to the local controller nodes of robots. In the following sections, we describe the functionality of these modules.
@@ III-B Frontier and reward management Server
This Section incorporates the central server in our previous work of [2] and will be briefly described here. The central server comprises of manager node and action servers that receive the list of frontier points from the frontier filter module. It quantifies them with the merged map information gain, replies with the next target goal position to the robot, and keeps track of already assigned goal points. As shown in Figure 1 manager node receives the filter point for each agent and acts as a communication gateway between the server and robot. The merge points action server creates a unique list of frontier points as MAX_PTS to be used by all the agents alongside their information gain value depending in the radius RAD and percentage of unknown cells inside it as PRC_UNK. The choose goals action server chooses the best goal position for each agent by considering the distance to previously assigned goal points as DIST_THRES and depending on the reward matrix computation.
@@ III-C Re-localization method
In AC-SLAM the scalar mapping of is used to quantify the uncertainty by reasoning over its Eigenvalues and determinant as described in [18]. As the robot explores the environment its uncertainly increases and without loop closure or help from neighboring robots, it loses its SLAM estimate. We quantify the uncertainty (D-Optimality [18]) as edge D-Optimality (D-Opti) as defined in Equation 2 where are the Eigenvalues of and is the dimension of each edge covariance matrix.
@@ IV SIMULATION RESULTS
The simulations111https://www.youtube.com/watch?v=6j3VBdnVcO8.. were carried out on ROS Noetic, Ubunto 20.04LTS on Intel Core i7®, with a system RAM of 32GB and NVIDIA RTX 1000 GPU. As described earlier, we modified the approach of [1] to multi-robot and implemented the proposed approach as mentioned in Section III using ORB SLAM2 [13] backend, RosBot222https://husarion.com/. equipped with RGBD sensor, LiDAR, and using Timed Elastic Band (TEB) and A-Star as local and global planners from the ROS navigation stack 333http://wiki.ros.org/navigation.. We used open-source modified House Environment (H.E) and Warehouse Environment (W.E) from Gazebo simulator444https://github.com/aws-robotics. measuring 157 and 260 respectively. Figure 2 shows the Gazebo images masked on top of the resulting O.G maps from UAV by flying it at 5 height in a lawn mover pattern and the final map of H.E. From the O.G maps in Figure 2a and 2b we can observe a difference in obstacle area as both environments have different nature of obstacles contributing to more ORB features for W.E. Figure 2c shows the resulting merged map from Robot_0, Robot_1 and UAV, and indicating the initial, final robot positions alongside their filtered frontiers. The ground truth O.G maps for local and global planners were generated using the Gazebo 2D Map plugin555https://github.com/marinaKollmitz/gazebo. which uses wavefront exploration.

### v2

@@ I INTRODUCTION
Current research is increasingly focusing on Active Collaborative SLAM (AC-SLAM), which leverages multiple robots working in unison. This collaborative method offers significant advantages, such as faster terrain mapping and robust operation in dynamic scenarios. However, the deployment of multiple robots introduces challenges like coordination, resource management, and sensor data integration. Moreover, merging individual robot contributions into a unified map poses complex computational and algorithmic challenges. Here, we propose an implementation of a visual AC-SLAM algorithm using heterogeneous agents and extend the work in [1] to a multi-agent system, where a team of ground robots and aerial robots (UAVs) collaboratively map and explore the environment. To achieve this aim, we propose a method that exploits ORB [2] feature information from different viewpoints of UAVs and ground robots into a common merged map, we detect frontiers in local and overlapped ground robot maps and propose a frontier filtering method that encourages the spread of agents and exploration of the environment. Additionally, we also present a re-localization scheme to guide robots to previously visited goal positions to favor loop closure.
@@ II-A Active Collaborative SLAM
[8] presents a decentralized method for a long-planning horizon of actions for exploration and maintains estimation uncertainties at a certain threshold. The active path planner uses a modified version of RRT* and an action is chosen that best minimizes the entropy change per distance traveled. The main advantage of this approach is that it maintains good pose estimation and encourages loop-closure trajectories. An interesting solution is given by a similar approach to the method proposed by [9] using a Relative Entropy (RE) optimization method which integrates motion planning with robot localization and selects trajectories that minimize the localization error and associated uncertainty bound. A planning-cost function is computed, which includes the uncertainty in the state in addition to the state and control cost.
@@ II-B Frontiers-based exploration
Frontiers play a pivotal role in augmenting the precision of robot localization by enabling intelligent exploration and data acquisition strategies, effectively reducing uncertainty, and enhancing the map building and localization processes. In [10] each frontier is segmented, a trajectory is planned for each segment, and the trajectory with the highest map segment covariance is selected from the global cost map. The work presented in [4] uses frontier exploration for autonomous exploration a utility function based on Shannon’s and Renyi entropy is used for the computation of the utility of paths. The method described by [11] uses a cost function that is somewhat similar to [5], it takes into consideration the discovery of the target area of a robot by another member of the swarm and switches from a frontier to a distance based navigation function to guide the robot toward the goal frontier. Frontiers-based coverage approaches in [12] divide the perception task into a broad exploration layer and a detailed mapping layer, making use of heterogeneous robots to carry out the two tasks while solving a Fixed Start Open Traveling Salesman Problem (FSOTSP). Once a frontier has been identified, the robot can use path planning algorithms to reach it and maximize the exploration while minimizing its SLAM uncertainly. In [13] the authors present decentralized autonomous exploration method which takes into account frontiers detected by other agent and inter robot exchange of frontiers. This method uses distance and information gain matrices for frontier utility computation.
@@ III METHODOLOGY
While many research works have been focused on collaborative strategies for SLAM, or single-robot active-SLAM, only a few works have dealt with AC-SLAM. However, these approaches present common limitations: a) they have high computational costs associated with the number of frontiers processed. b) They fail to encourage the spread of robots into the environment by not taking into account the common frontier points from the agents. c) The uncertainty is quantified by a scalar mapping of the entire pose graph covariance matrix which may become very large, especially in landmark-based SLAM methods increasing the computational cost. d) In case of SLAM failure or high uncertainty, there is no re-localization method. Furthermore, they do not explicitly implement strategies for efficient management of frontiers to speed up map discovery and robot localization. In this work, we propose a visual AC-SLAM approach that deals with overcoming these limitations using a fleet of ground and aerial robots and exploiting visual features from different viewpoints of both heterogeneous robots. Our proposed method outlines a strategy aimed at reducing the number of frontiers for reward computation and distributing robots within the environment, thereby facilitating exploration and mapping. Furthermore, we also present a re-localization method quantifying the SLAM pose graph edge covariance matrix hereby allowing the robots to visit already visited goal positions for re-localization and uncertainty reduction for a better SLAM estimate. We leverage a combination of reward metrics, common map points, and merged map information gain to refine the goal selection.
@@
We expand the method of [1] to a multi-robot AC-SLAM and incorporate the utility function from our previous work in [15] which considers frontier path entropy to exploit map uncertainty for utility computation. Utilizing the Intersection over Union (IoU) overlapping map information, the number of frontiers is significantly reduced for efficient frontier sharing and exploration. Utilizing edge D-Optimality as an uncertainty metric, we propose a re-localization scheme to encourage loop closure and reduce uncertainty. In Figure 1 we show the resultant architecture of the proposed system implemented in ROS Noetic using actionlib111http://wiki.ros.org/actionlib library for UAV and ground robots (Robot_0… Robot_N). Each robot/UAV runs its own instance of ORBSLAM2 using onboard RGBD sensors. The resultant ORB features are converted to a 3D voxel map (Octomap) [16] for efficient representation using Octo and Grid Mapper module222http://wiki.ros.org/octomap_server. For each of the ground robot and drone, we limit the maximum and minimum height of the Octomap to 0.5 and 1 meters () respectively, thus limiting the ground robots to ignore ORB features above 0.5 high and for drones to ignore them of less than 1. We then project this Octomap to a 2D Occupancy Grid (O.G) map and apply frontier detection methods using OpenCV and RRT-exploration from [1] to detect local frontiers in each robot’s local map. The MapMerger module merges the offline drone O.G maps with those of ground robots with world reference frame set to that of Robot_0. The IoU mapper computes the overlapped O.G map between two robots and detects frontiers and penalizes each robot’s local frontiers thus reducing them. Section III-A elaborates this process. Once the frontiers have been filtered they are passed to the frontier management and reward computation server as discussed in Section III-B which performs further reduction and updates the rewards before finally assigning goal frontiers to the local controller nodes of robots. In the following sections, we describe the functionality of these modules.
@@ III-B Frontier and reward management Server
This section incorporates the central server in our previous work of [3] and will be briefly described here. The central server comprises of manager node and action servers that receive the list of frontier points from the frontier filter module. It quantifies them with the merged map information gain and returns the next target goal point to the robot, it also keeps track of already assigned goal points. As shown in Figure 1 manager node receives the filtered points from each agent and acts as a communication gateway between the server and robot. The merge points action server then creates a unique list of frontier points (which is less than the maximum points MAX_PTS) to be used by all the agents alongside their information gain value depending on the radius RAD and percentage of unknown cells inside it as PRC_UNK. The choose goals action server determines the best goal position for each agent by considering the distance to previously assigned goal points as DIST_THRES and depending on the reward matrix computation.
@@ III-C Re-localization method
The scalar mapping of is used to quantify the uncertainty by reasoning over its Eigenvalues and determinant as described in [18]. As the robot explores the environment its uncertainly increases and without loop closure or localization help from neighboring robots, it loses its SLAM estimate. We quantify the uncertainty (D-Optimality [18]) as edge D-Optimality (D-Opti) as defined in Equation 2 where are the Eigenvalues of and is the dimension of each edge covariance matrix.
@@ IV SIMULATION RESULTS
The simulations111https://www.youtube.com/watch?v=6j3VBdnVcO8.. were carried out on ROS Noetic, Ubunto 20.04LTS on Intel Core i7®, with a system RAM of 32GB and NVIDIA RTX 1000 GPU. As described earlier, we modified the approach of [1] to multi-robot and implemented the proposed approach as mentioned in Section III using ORB SLAM2 [14] backend, RosBot222https://husarion.com/. equipped with RGBD sensor, LiDAR, and using Timed Elastic Band (TEB) and A∗ as local and global planners respectively from the ROS navigation stack 333http://wiki.ros.org/navigation.. We used open-source modified House Environment (H.E) and Warehouse Environment (W.E) from Gazebo simulator444https://github.com/aws-robotics. measuring 157 and 260 respectively. Figure 2 shows the Gazebo images masked on top of the resulting O.G maps from UAV by flying it at 5 height in a lawn mover pattern and the final map of H.E. From the O.G maps in Figure 2a and 2b we can observe a difference in obstacle area as both environments have different nature of obstacles contributing to more ORB features for W.E. Figure 2c shows the resulting merged map from Robot_0, Robot_1 and UAV, and indicating the initial, final robot positions alongside their filtered frontiers. The ground truth O.G maps for local and global planners were generated using the Gazebo 2D Map plugin555https://github.com/marinaKollmitz/gazebo. which uses wavefront exploration.